OpenStudy (anonymous):

find the slope of a line parrallel to 5x + 2y = 6.

OpenStudy (arabpride):

turn into y = mx + b form ~then check to see of the lines contain the same slope to determine whether it is a parallel line or not.

OpenStudy (anonymous):

m=−5/2 is the slope of that line.

OpenStudy (anonymous):

so thats your answer

OpenStudy (arabpride):

5x + 2y = 6. -5x -5x 2y = -5x + 6 /2 /2 y = -5/2x + 3 now check to see if both lines contain the slope -5/2x ~ if so, they are parallel

OpenStudy (anonymous):

turn into y=mx+b so it looks like y=-5/2x+3 so your slope is m=-5/2 which would be the same for a parallel line.

OpenStudy (anonymous):

thank you!

OpenStudy (arabpride):

Welcome *^.^*